You are among 237 citizens who have been selected for jury duty. Using this base, determine the following probabilities: a. That you will be one of 45 citizens selected for a jury panel b. That you will then be one of 12 citizens selected to serve as members of the jury An expert already answered this question but I did not understand I need more details.
Answer #1

Answer:

Given,

Total number of citizens = 237

a)

To give the probability that you will be one of 45 citizens selected for a jury panel

Required probability = 45C1 / 237C1 [ here nCr = n!/(n-r)!*r!]

= 45 / 237

= 0.1899

Required probability = 0.1899

b)

To give the probability that you will then be one of 12 citizens selected to serve as members of the jury

i.e.,

= 12C1 / 237C1

= 12 / 237

= 0.0506

Required probability = 0.0506